OpenStudy (anonymous):

passing through:(4,-3) & perpendicular to y=-3/2x+2

OpenStudy (anonymous):

y=2/3x+b -3=8/3+b b=-17/3 y=(2/3)x-17/3

OpenStudy (anonymous):

Thank You